A card is drawn from pack of 52 cards and it is thrown away.Then another card is drawn.Find the probability that it is a queen.
1
Expert's answer
2012-10-09T08:42:35-0400
According to the full probability conception, we have:

P(second card is a queen) = P(to draw a queen if the first card was not a queen) +
& + P(to draw a queen if the first card was a queen).

We know that

P(to draw a queen if the first card was not a queen) = 4/53,
P(to draw a queen if the first card was a queen) = 3/53.

So,

P(second card is a queen) = 4/53 + 3/53 = 7/53.